OpenStudy (anonymous):

The sum of FOUR consecutive odd integers is 208. What are the numbers?

OpenStudy (anonymous):

I used a shortcut.... XP 208/4 =52 <== this gives the average of the 4 consecutive odd #s so the numbers are 49, 51, 53, 55 You can also use algebra...... Let x = first number x+2 = 2nd number x+4 = 3rd number x + 6 = 4th number x+x+2+x+4+x+6=208 and solve for x.... etc

OpenStudy (anonymous):

thnx :)

OpenStudy (anonymous):

this is so awesome thanks for the new outlook :) Your so smart

OpenStudy (anonymous):

lol Thanks, and you're welcome.